Find the unknown constants, a1 and a2 in the partial fraction expansion 3 s + 15 ( s + 6 ) 2 = a 1 ( s + 6 ) 2 + a 2 ( s + 6 ) .
Answer #1

92 38 +15 4 F (8+6)2 (8+62 (+) 9,+ 38+15 = 2 (8+6) = > 38+15= (ait 692) + ais Comparing both sides 9, +692= 15 92 = 3 ait 6x3